a father is 42 years old,his son is 10 years old and his daughter is 16 years old.
a) In how many years will the age of the father be the triple of that of his son?
b) In how many years will the age of the father be the sum of those of his children?

Found 2 solutions by lwsshak3, hamsanash1981@gmail.com:
Answer by lwsshak3(11628)

a father is 42 years old,his son is 10 years old and his daughter is 16 years old.
a) In how many years will the age of the father be the triple of that of his son?
b) In how many years will the age of the father be the sum of those of his children?
***
let x=years
a) (32+x)/(10+x)=3
32-x=30-3x
2x=12
x=6
In how many years will the age of the father be the triple of that of his son? 6
..
b)(32+x)=(10+x+16+x)
x=6
In how many years will the age of the father be the sum of those of his children? 6

Answer by hamsanash1981@gmail.com(151)

a) let be the no of years taken to triple the age of his son.
then father's age = 42+x , son's age = 10+x.
42+x = 3(10+x)
42 + x = 30 + 3x
12 = 2x
x = 6
b) Let x be the number of years. father's age = 42+x , son's age = 10+x, daughter's age = 16+x
42+x = 10+x + 16+ x
x=16